OpenStudy (wcrmelissa2001):

help please :) question on graphical transformations. 1. When transforming asymptotes, (i.e.1/x) and there's a negative sign, when do we reflect in x axis and when do we reflect in y axis? 2. If I have y=3+x/2 and i reflect it in x axis, does it become 3-x/2 or -(3+x/2)? Thank you :)

OpenStudy (welshfella):

https://www.desmos.com/calculator/ybgvttqlmo I'm not sure i understand your question but the above shows graphs of y = 1/x and y = -1/x

OpenStudy (wcrmelissa2001):

sorry which question don't you understand?

OpenStudy (wcrmelissa2001):

oh basically for asymptotes, when you reflect them you need to know if it's reflected in the x axis or y axis. And you can tell this by where the minus signs are located. So if i'm reflecting in y axis, where should the minus sign be? (and same goes for x axis)

OpenStudy (wcrmelissa2001):

the above graph can be a reflection in y or x axis?

OpenStudy (welshfella):

No waitI take a look at that graph is it a reflection in the x-axis ? . I think it is

OpenStudy (wcrmelissa2001):

ohhhhh

OpenStudy (welshfella):

Yes it is

OpenStudy (wcrmelissa2001):

yeah. i see it! but when would it be a reflection in the y axis?

OpenStudy (welshfella):

i get confused with these sometimes. Yea it is.

OpenStudy (wcrmelissa2001):

but when would it be a reflection in the y axis?

OpenStudy (welshfella):

You can also take it to be a reflection across the y-axis because the 2 graphs are symmetrical about the origin.

OpenStudy (welshfella):

agreed?

OpenStudy (welshfella):

do you see that?

OpenStudy (welshfella):

or am i wrong

OpenStudy (welshfella):

as for the second part when you reflect across the x axis the y values become negative

OpenStudy (wcrmelissa2001):

wait i thought this graph can't be a reflection in y axis?

OpenStudy (welshfella):

why not?

OpenStudy (wcrmelissa2001):

no wait. i just confused myself

OpenStudy (wcrmelissa2001):

wait let's try plotting 1/x-5

OpenStudy (welshfella):

https://www.desmos.com/calculator/0xwnolr5rq there's the second part

OpenStudy (wcrmelissa2001):

no wait it's the same what am i doing

OpenStudy (wcrmelissa2001):

thanks

OpenStudy (welshfella):

- your second choice was correct in Q 2.

OpenStudy (welshfella):

yw

OpenStudy (wcrmelissa2001):

okie ;D for first part right does this graph help me in any way https://www.desmos.com/calculator/ybgvttqlmo

OpenStudy (wcrmelissa2001):

this is the y xis one right

OpenStudy (welshfella):

well you can look upon it as a reflection in the x axis or e reflection in y-axis.

OpenStudy (welshfella):

both give the same image.

OpenStudy (welshfella):

https://www.desmos.com/calculator/odu9pg9z6y here are the graphs of 1 / (x - 5) and y = -1/(x -5) this is a reflection in the x axis but not in the y axis 9 it can also be regarded as a reflection in the line x = 5))

OpenStudy (welshfella):

the question is what would be a reflection in the y axis for this function ?

OpenStudy (wcrmelissa2001):

no. just when i need to reflect it in the y axis if its an asymptote

OpenStudy (welshfella):

well i guess we change all x values to negative

OpenStudy (wcrmelissa2001):

hmm ok thanks!

OpenStudy (welshfella):

https://www.desmos.com/calculator/odu9pg9z6y check that out

OpenStudy (wcrmelissa2001):

woah. clears things up more! thanks man

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

hmm i didnt mean to send that one this is the reflection in y axis

OpenStudy (wcrmelissa2001):

haha it's ok. i think i get it

OpenStudy (welshfella):

https://www.desmos.com/calculator/var6b5xd50

OpenStudy (wcrmelissa2001):

THANKS :D

OpenStudy (welshfella):

yw